# Asymptotic and compound e-values: multiple testing and empirical Bayes

### v2

@@ 7.2 The best simple separable compound e-values
$\sum_{k\in\mathcal{K}}\mathbb{E}^{\mathbb{P}_{k}}[s(X_{k})]=K.$
$\int s(x)\overline{p}^{\circ}(x)\mathrm{d}\nu(x)=1.$
7.3 Compound universal inference
We now slightly modify the setting of Section 7.2 to accommodate composite null hypotheses with nuisance parameters. We continue to suppose that $X=(X_{k}:k\in\mathcal{K})$, where $X_{k}$ takes values in a space $\mathcal{X}$ and that the $k$-th hypothesis $H_{k}$ is specified solely in terms of the marginal distribution of $X_{k}$. However, we now suppose that $H_{k}$ is composite and that $k$-th null marginal may be parameterized by a nuisance parameter $\psi_{k}$ that lies in a parameter space $\Psi$. Concretely, $\mathcal{P}_{k}=\{\mathbb{P}:\mathrm{d}\mathbb{P}_{k}/\mathrm{d}\nu=p^{\circ}_{\psi_{k}}\text{ for some }\psi_{k}\in\Psi\}$, where $\{p^{\circ}_{\psi}:\psi\in\Psi\}$ is a family of $\mathrm{d}\nu$-densities on $\mathcal{X}$. We seek to test each composite null $H_{k}$ against the alternative $\{\mathbb{Q}:\mathrm{d}\mathbb{Q}_{k}/\mathrm{d}\nu=q^{\circ}\}$, where $q^{\circ}$ is a further $\mathrm{d}\nu$-density on $\mathcal{X}$ (which we choose to be the same for all $k\in\mathcal{K}$).
In the above setting, universal inference (UI) of Wasserman et al. (2020) provides a flexible and practical way of constructing separable e-values: [^6] Throughout this section we assume that infima are attained.
If we instead seek to construct compound e-values, we propose compound universal inference (CUI) in which we profile over the nuisance parameters appearing in the optimal separable compound e-values of Theorem 7.2:
$E_{k}^{\mathrm{CUI}}=\frac{Kq^{\circ}(X_{k})}{\sum_{j\in\mathcal{K}}p^{\circ}_{\widehat{\psi}_{j}}(X_{k})}\,\text{ for all }\,k\in\mathcal{K},\,\,\,$ (17)
$\text{where}\;\widehat{\psi}_{1},\dots,\widehat{\psi}_{K}\in\operatorname*{argmin}\left\{\sum_{k\in\mathcal{K}}\frac{Kq^{\circ}(X_{k})}{\sum_{j\in\mathcal{K}}p^{\circ}_{\psi_{j}}(X_{k})}\,:\,\psi_{1},\dotsc,\psi_{K}\in\Psi\right\}.$
We also consider a variation of the above that we call empirical Bayes universal inference (EBUI):777 Nguyen and Gupta (2023) use universal inference for an empirical Bayes application. By contrast, here we use ideas from empirical Bayes and universal inference to provide a new construction of compound e-values.
$E_{k}^{\mathrm{EBUI}}=\frac{q^{\circ}(X_{k})}{\int p^{\circ}_{\psi}(X_{k})\mathrm{d}\widehat{G}(\psi)}\,\text{ for all }\,k\in\mathcal{K},\,\,\,$ (18)
$\text{where}\;\widehat{G}\in\operatorname*{argmin}\left\{\sum_{k\in\mathcal{K}}\frac{q^{\circ}(X_{k})}{\int p^{\circ}_{\psi}(X_{k})\mathrm{d}G(\psi)}\,:\,G\text{ distribution supported on }\Psi\right\}.$
We have the following result according to which compound universal inference leads to compound e-values whose sum is at least as large as the universal inference e-values. This highlights potential power gains from considering compound e-values instead of e-values.
@@ 7.3 Compound universal inference
All of the following are compound e-variables for $\mathcal{P}_{1},\ldots,\mathcal{P}_{K}$:
  • $E_{1}^{\mathrm{UI}},\dots,E_{K}^{\mathrm{UI}}$,
  • $E_{1}^{\mathrm{CUI}},\dots,E_{K}^{\mathrm{CUI}}$,
  • $E_{1}^{\mathrm{EBUI}},\dots,E_{K}^{\mathrm{EBUI}}$.
Moreover it holds that:
$\sum_{k\in\mathcal{K}}E_{k}^{\mathrm{CUI}}\geq\sum_{k\in\mathcal{K}}E_{k}^{\mathrm{EBUI}}\geq\sum_{k\in\mathcal{K}}E_{k}^{\mathrm{UI}}.$
Proof.
We first prove the inequalities. Let $\widehat{\psi}_{1},\dots,\widehat{\psi}_{K}$ be the minimizers in (17). Then, the distribution $\sum_{k\in\mathcal{K}}\delta_{\widehat{\psi}_{k}}/K$ is feasible for the minimization in (18) and leads to the same objective value, proving the first inequality. Next:
$\sum_{k\in\mathcal{K}}E_{k}^{\mathrm{EBUI}}$ $=\inf_{G}\sum_{k\in\mathcal{K}}\frac{q^{\circ}(X_{k})}{\int p^{\circ}_{\psi}(X_{k})\mathrm{d}G(\psi)}\geq\sum_{k\in\mathcal{K}}\inf_{G}\frac{q^{\circ}(X_{k})}{\int p^{\circ}_{\psi}(X_{k})\mathrm{d}G(\psi)}=\sum_{k\in\mathcal{K}}\inf_{\psi_{k}\in\Psi}\frac{q^{\circ}(X_{k})}{p^{\circ}_{\psi_{k}}(X_{k})}=\sum_{k\in\mathcal{K}}E_{k}^{\mathrm{UI}}.$
Now, take any $\mathbb{P}\in\mathcal{P}$. For any $k$ with $\mathbb{P}\in\mathcal{P}_{k}$, let $\psi_{k}$ be such that $\mathrm{d}\mathbb{P}_{k}/\mathrm{d}\nu=p^{\circ}_{\psi_{k}}$ (which exists by the definition of $\mathcal{P}_{k}$ in this section). For all other $k$, let $\psi_{k}\in\Psi$ be arbitrary. Then by construction,
$\sum_{k\in\mathcal{K}}\frac{Kq^{\circ}(X_{k})}{\sum_{j\in\mathcal{K}}p^{\circ}_{\psi_{j}}(X_{k})}\geq\sum_{k\in\mathcal{K}}E_{k}^{\mathrm{CUI}}.$
Arguing as in the proof of Theorem 7.2, the summands in the left hand side of the above inequality are compound e-values. Hence $E_{1}^{\mathrm{CUI}},\dots,E_{K}^{\mathrm{CUI}}$ are compound e-values (and similarly for EBUI and UI). ∎

### v3

@@ 6.1.1 Localized compound universal inference for the simultaneous t-test
This follows by the Dvoretzky–Kiefer–Wolfowitz inequality with Massart’s tight constant (Massart, 1990) and Bretagnolle’s argument (Bretagnolle, 1981); see also Donoho and Jin (2006, Lemma 7.1). The additive factor of $1$ on the right-hand side is needed because $\hat{\sigma}_{k}^{2}$ are not identically distributed. ∎
The construction is akin to the $F$-Localization of Ignatiadis and Wager (2022) with the difference that we do not require that $\sigma_{k}^{2}\,{\buildrel\text{iid}\over{\sim}\,}G$ and instead construct the localization in the compound setting.
Remark 6.2 (Computation of CUI.).
Computing CUI amounts to solving the optimization problem in (13). To do this in practice, we proceed as follows. First, following Ignatiadis and Wager (2022), we discretize $\mathcal{G}$ by taking a covering of $[\underaccent{\bar}{\sigma}^{2},\bar{\sigma}^{2}]$ with $B$ points $\tilde{\sigma}_{1}^{2},\dots,\tilde{\sigma}_{B}^{2}\in[\underaccent{\bar}{\sigma}^{2},\bar{\sigma}^{2}]$ and considering all distributions of the form $G=\sum_{\ell=1}^{B}\pi_{\ell}\delta_{\tilde{\sigma}_{\ell}^{2}}$, where $(\pi_{1},\dots,\pi_{B})$ lies on the probability simplex. Second, we relax the Kolmogorov-Smirnov constraint in (19) by enforcing it only on a finite subset $\mathcal{T}\subseteq(0,\infty)$; e.g., we can let $\mathcal{T}$ be a subset of the realized sample variances $\hat{\sigma}_{1}^{2},\dotsc,\hat{\sigma}_{K}^{2}$. This second discretization step is always conservative, that is, it makes the localization strictly larger. Our computational approximation to $E_{k}^{\mathrm{CUI}}$ is then given by $q^{\circ}(X_{k})/\mathrm{obj}$, where $\mathrm{obj}$ is the optimal value of the following linear program,
$\underset{\pi_{1},\ldots,\pi_{B}}{\text{maximize}}\quad$ $\sum_{\ell=1}^{B}\pi_{\ell}p^{\circ}_{\tilde{\sigma}_{\ell}^{2}}(X_{k})$
subject to $\sum_{\ell=1}^{B}\pi_{\ell}=1,\quad\pi_{\ell}\geq 0,\quad\ell=1,\ldots,B,$
$\left|\frac{1}{K}\sum_{j\in\mathcal{K}}\mathds{1}_{\{\hat{\sigma}_{j}^{2}\leq t\}}-\sum_{\ell=1}^{B}\pi_{\ell}\int_{0}^{t}p(u\mid\tilde{\sigma}_{\ell}^{2})\,\mathrm{d}u\right|\leq\sqrt{\frac{1+\log(2/\delta)}{2K}},\quad t\in\mathcal{T}.$
6.1.2 Optimal discovery compound e-values via empirical Bayes
In this section, we take a more direct approach to approximating the optimal discovery compound e-values. We directly use empirical Bayes to estimate the compound empirical distribution $G(\boldsymbol{\sigma}^{2})$ in (18).
It will be convenient to parameterize the distribution of $X_{k}$ by $(\lambda_{k},\sigma_{k}^{2})$ where $\lambda_{k}=\mu_{k}/\sigma_{k}$, that is, for $x\in\mathbb{R}^{n}$, we write:
$p_{\lambda,\sigma^{2}}(x)=\frac{1}{(2\pi\sigma^{2})^{n/2}}\prod_{i=1}^{n}\exp\left(-\frac{(x_{i}-\lambda\sigma)^{2}}{2\sigma^{2}}\right),\;\;\;p_{\sigma^{2}}(x)=p_{0,\sigma^{2}}(x).$
In this way, given a measure $Q$ over ×$\mathbb{R}\times(0,\infty)$ and $G$ over $(0,\infty)$, we may define the Bayes factor
$E(x;G,Q)=\frac{\int p_{\lambda,\sigma^{2}}(x)\mathrm{d}Q(\lambda,\sigma^{2})}{\int p_{\sigma^{2}}(x)\mathrm{d}G(\sigma^{2})}.$ (20)
For fixed $Q$ and with $G=G(\boldsymbol{\sigma}^{2})$, it holds that
$E(x;G(\boldsymbol{\sigma}^{2}),Q)=\frac{K\int p_{\lambda,\sigma^{2}}(x)\mathrm{d}Q(\lambda,\sigma^{2})}{\sum_{j\in\mathcal{K}}p_{\sigma_{j}^{2}}(x)},$
and so, $E(X_{k};G(\boldsymbol{\sigma}^{2}),Q)$, $k\in\mathcal{K}$, are optimal discovery compound e-variables. We propose to construct asymptotic compound e-variables
$E_{k}=E(X_{k};\widehat{G},\widehat{Q}),$ (21)
by plugging in specific estimates $\widehat{G}$ of $G$ and $\widehat{Q}$ of $Q$ as described below. We estimate $\widehat{G}$ based on the nonparametric maximum likelihood estimator (NPMLE) of $G$ by Robbins (1950) and Kiefer and Wolfowitz (1956) based on $\hat{\sigma}_{1}^{2},\ldots,\hat{\sigma}_{K}^{2}$:
$\widehat{G}\in\operatorname*{argmax}\left\{\sum_{k\in\mathcal{K}}\log\left(f_{G}(\hat{\sigma}_{K}^{2})\right)\,:\,G\in\mathcal{G}\right\},\;\;$ (22)
where we recall that $\mathcal{G}$ is the class of all distributions supported on $[\underaccent{\bar}{\sigma}^{2},\bar{\sigma}^{2}]$. Moreover we let $\widehat{Q}=\widehat{H}\otimes\widehat{G}$ where $\widehat{H}$ is a distribution on $\mathbb{R}$ estimated based on $X_{1},\ldots,X_{K}$ and that is supported on $[\underaccent{\bar}{\lambda},\bar{\lambda}]\subseteq(0,\infty)$, where $\underaccent{\bar}{\lambda},\bar{\lambda}$ are fixed.
Theorem 6.3.
Consider asymptotics with $n\geq 3$ fixed and $K\to\infty$. Assume there exists $H^{*}$ such that $\widehat{H}\stackrel{{\scriptstyle d}}{{\rightsquigarrow}}H^{*}$ almost surely. Then $E_{1},\ldots,E_{K}$ in (21) are strongly asymptotic compound e-variables for $\mathcal{P}_{1},\ldots,\mathcal{P}_{K}$ under $\mathcal{P}^{\mathrm{N}}$.
Theorem 6.3 is a purely frequentist result. As is common in compound decision theory, we “pretend” that $\sigma_{k}^{2}\sim G$ when computing the NPMLE, however, in the theorem statements, $\sigma_{1}^{2},\ldots,\sigma_{K}^{2}$ are fixed (deterministic). Moreover, $H^{*}$ can be any distribution and does not need to represent the data-generating distribution.
Since the ODP was first described by Storey (2007), several authors have proposed data-driven implementations of the ODP when the nuisance parameters are unknown including Storey et al. (2007); Guindani et al. (2009); Woo et al. (2011). The approach closest to ours is the one of Noma and Matsui (2012, 2013) who also pursue an empirical Bayes approach; however, they only consider parametric specifications for $G$ (while allowing for nonparametric $H$). To the best of our knowledge, Theorem 6.3 presents the first theoretical guarantee for data-driven ODP and together with Theorem 4.4 it implies that we can use the empirical Bayes optimal discovery compound e-values alongside the e-BH procedure to asymptotically control the FDR.
Remark 6.4 (Computation of NPMLE).
In practice we compute the NPMLE (22) using the standard strategy proposed by Koenker and Mizera (2014): we first discretize $G$, as described in Remark 6.2, and this discretization reduces (22) to a finite conic programming problem that can be solved by interior point methods.
6.1.3 Simulation study
We demonstrate the improved power of our two constructions through a simulation study. We simulate from the normal t-test model with $K=2000$ and let $n\in\{5,10\}$ be a simulation parameter. We consider two settings for $\sigma_{k}^{2}$, either all $\sigma_{k}^{2}=1$ or $\sigma_{k}^{2}\,{\buildrel\text{iid}\over{\sim}\,}\mathrm{Unif}[0.5,2]$. Moreover we set $\lambda_{k}=0$ for $1800$ hypotheses (nulls), and $\lambda_{k}=\xi/\sqrt{n}$ for the other 200 hypotheses. We call $\xi$ the effect size and vary it in the interval $[2,6]$ in our simulations.
We apply e-BH at level $\alpha=0.1$ and compare the following methods. We provide all methods with access to $\xi/\sqrt{n}$ and they all thus set $H=\delta_{\xi/\sqrt{n}}$. This is favorable for all methods but allows us to isolate the impact of the denominators in constructing (asymptotic, compound) e-values.
  1. 1.
    z-Oracle: This is an idealized method that has access to the true $\sigma_{k}^{2}$ for all $k$ and sets $E_{k}^{z}=E(X_{k};\delta_{\sigma_{k}^{2}},H\otimes\delta_{\sigma_{k}^{2}})$.
  2. 2.
    EB-oracle: This is another idealized method with access to the empirical distribution $G(\boldsymbol{\sigma}^{2})$. It is equal to the (oracle) optimal discovery compound e-values $E(X_{k};G(\boldsymbol{\sigma}^{2}),H\otimes G(\boldsymbol{\sigma}^{2}))$.
For the remaining methods, we use the universe in (17) with $\underaccent{\bar}{\sigma}^{2}=10^{-3}$ and $\bar{\sigma}^{2}=10^{3}$. The remaining methods are data-driven (i.e., are not given any extra information about the nuisance parameters beyond the wide bounds $\sigma_{k}^{2}\in[\underaccent{\bar}{\sigma}^{2},\bar{\sigma}^{2}]$). For the discretization of $G$ (described in Remarks 6.2 and 6.4), we take $B=600$ and consider a logarithmically equispaced grid from $\underaccent{\bar}{\sigma}^{2}$ to $\bar{\sigma}^{2}$. Below, $\widehat{G}$ refers to the NPMLE.
  1. 3.
    t-test: This is the scale invariant e-value studied in depth by Wang and Ramdas (2025). The $k$-th e-value is a measurable function of the t-statistic $T_{k}=\sqrt{n}\bar{X}_{k}/\hat{\sigma}_{k}$.
  2. 4.
    Empirical Bayes (EB): The plug-in compound e-values studied in Theorem 6.3, that is, $E_{k}^{\text{EB}}=E(X_{k};\widehat{G},H\otimes\widehat{G})$.
  3. 5.
    Universal inference (UI): As in (11) but with numerator given by $\int p_{\lambda,\sigma^{2}}(x)\mathrm{d}(H\otimes\widehat{G})(\lambda,\sigma^{2})$.
  4. 6.
    Compound universal inference (CUI): as in Section 6.1.1 with the same numerator as for UI and the localization in (19) with $\delta=0.01$. We apply e-BH at level $0.09$ rather than $0.1$. [^7] Formally, UI and CUI are not fully justified as they estimate the numerator based on the data. However, we expect them to be more conservative than EB, and so include them as a comparison of different treatments for the denominator.
We evaluate results in terms of FDR and Power, which we define as the expected proportion of true discoveries among non-null hypotheses. These metrics are estimated by averaging the results over 200 Monte Carlo replications of each simulation.
Figure 1: Power comparison of six multiple testing procedures under different sample sizes ($n\in\{5,10\}$) and variance settings ($\sigma_{k}^{2}=1$ vs. $\sigma_{k}^{2}\sim\mathrm{Unif}[0.5,2]$). Methods include Z-test (Oracle), Separable (Oracle), EB (Empirical Bayes), CUI (Compound Universal Inference), t-test, and UI (Universal Inference). Among data-driven methods (i.e., methods that do not have oracle knowledge regarding the variances), our proposed EB method has by far the largest power. Moreover, EB is indistinguishable from its oracle counterpart (EB-Oracle).
The results of the simulation are shown in Figure 1. All methods control FDR at the nominal $\alpha=0.1$ (not shown). In all cases, the most powerful methods is z-Oracle, which has the most information. When all $\sigma_{k}^{2}=1$, then z-Oracle and EB-Oracle are identical (since knowledge that $G(\boldsymbol{\sigma}^{2})=\delta_{1}$ implies that $\sigma_{k}^{2}=1$ for all $k$). Meanwhile, our data-driven version (EB), is effectively indistinguishable from EB-oracle, showing that the NPMLE effectively learns $G(\boldsymbol{\sigma}^{2})$. CUI is less powerful than EB, but has more power than standard e-value methods (UI and t). The relative performance of UI and t depends on the simulation, see Wang and Ramdas (2025) for further discussion. Overall, this simulation demonstrates the power benefits by using asymptotic compound e-values instead of (bona fide) e-values.
6.2 Asymptotic e-values and compound e-values beyond normality
In this section, we demonstrate how to construct asymptotic (compound) e-values without assuming normality.
6.2.1 A prototypical example of an asymptotic e-value
We first provide a prototypical construction of asymptotic e-variables along the lines of the central limit theorem (as discussed in Section 3.2). We consider the setting described in the beginning of Section 6 and also use the summary statistics in (16) (even though they are no longer sufficient in the absence of normal errors). We focus on the case of a single hypothesis, i.e., $K=1$ and consider asymptotics as $n\to\infty$. For this reason, we omit the subscript $k$, and write e.g., $\bar{X}^{(n)}=\bar{X}_{1}^{(n)}$. We seek to test the hypothesis $\mathcal{P}_{1}$ in (15).
For fixed $\lambda\in\mathbb{R}$, let
and
The next proposition justifies that (23) and (24) define sequences of strongly asymptotic e-variables.
Theorem 6.5.
The sequences of random variables $E^{(n)}$ in (23) and $\tilde{E}^{(n)}$ in (24) are strongly asymptotic e-variables for $\mathcal{P}_{1}$. If $\hat{\sigma}^{(n)}$ is used in place of $S^{(n)}$ in (23) and (24), then $E^{(n)}$ and $\tilde{E}^{(n)}$ are asymptotic e-variables for $\mathcal{P}_{1}$. In either case, for $\lambda>0$, $E^{(n)}$ (resp. $\tilde{E}^{(n)}$) grows to infinity with probability $1$ under the alternative that $X^{1},\dots,X^{n}$ are drawn iid from $\mathbb{Q}$ with $\mathbb{E}^{\mathbb{Q}}[X^{i}]>0$ (resp. $\mathbb{E}^{\mathbb{Q}}[X^{i}]\neq 0$) and $\mathrm{var}^{\mathbb{Q}}(X^{i})<\infty$.
is a strongly asymptotic e-variable.
One central purpose of defining asymptotic e-variables is to handle cases where one does not know how to construct a nonasymptotic e-variable. Indeed, for the above $\mathcal{P}_{1}$, we do not know of nonasymptotic e-variables (other than constants).
6.2.2 A prototypical construction of asymptotic compound e-values
We next describe a prototypical model of constructing asymptotic compound e-variables. The construction appeared in Ignatiadis et al. (2024) (assuming normality as in Section 6.1) and here we generalize the construction and show that it does not require normality. We start with the summary statistics in (16), and will consider asymptotics with $\max\{n,K\}\to\infty$. Ignatiadis et al. (2024) were interested in computing approximate compound e-variables that satisfy $E_{k}^{(n)}\propto(S_{k}^{(n)})^{2}$ for all $k\in\mathcal{K}$. To motivate the construction that follows, fix any $\mathbb{P}\in\mathcal{P}$ and write $\mu_{k}(\mathbb{P})=\mathbb{E}^{\mathbb{P}_{k}}[X_{k}^{i}]$ and $\sigma_{k}^{2}(\mathbb{P})=\mathrm{var}^{\mathbb{P}_{k}}(X_{k}^{i})$. Observe that for any $k\in\mathcal{K}$,
We will show that the above are asymptotic compound e-variables under an asymptotic setup indexed by $m\in\mathbb{N}$ such that $n=n(m)$, $K=K(m)$ and $\max\{n,K\}\to\infty$ as $m\to\infty$. Our formal result will be stated in terms of two universes of distributions. The first set imposes some restrictions on moments of the marginals,
$\mathcal{P}^{\mathrm{B}}=\{\mathbb{P}\in\mathcal{P}\;:\;\mathbb{E}^{\mathbb{P}_{k}}[|X_{k}^{i}|^{2(1+\delta)}]\leq C,\;\mathrm{var}^{\mathbb{P}_{k}}(X_{k}^{i})\geq c\;\text{ for all }\;k\},$
where $\delta\in(0,1],c,C\in(0,\infty)$ are constants. The second set is equal to the universe $\mathcal{P}^{\mathrm{N}}$ in (17).
@@ 6.2.2 A prototypical construction of asymptotic compound e-values
  1. (i)
  2. (ii)
7 An alternate definition of approximate compound p-values and e-values
In this section we provide an alternative definition in lieu of Definition 3.13 for approximate compound p-variables and e-variables. This definition has the advantage that it is easier to satisfy, but it provides weaker guarantees. We include it because the asymptotic notion it leads to for compound p-variables was central to the work of Ignatiadis and Sen (2025).
Definition 7.1 (*Approximate [^8] We use the asterisk (*approximate) to distinguish between two related but distinct definitions within this paper. We expect other authors may reasonably use “$(\varepsilon,\delta)$-approximate compound p-variables” or “$(\varepsilon,\delta)$-approximate compound e-variables” for either notion, provided the definition is clearly stated. compound p-variables and e-variables).
  1. (i)
    Nonnegative random variables $P_{1},\dots,P_{K}$ are $(\varepsilon,\delta)^{*}$-approximate compound p-variables for $(\mathcal{P}_{1},\dots,\mathcal{P}_{K})$ under $\mathcal{P}$ if
  2. (ii)
    $[0,\infty]$-valued random variables $E_{1},\dots,E_{K}$ are $(\varepsilon,\delta)^{*}$-approximate compound e-variables for $(\mathcal{P}_{1},\dots,\mathcal{P}_{K})$ under $\mathcal{P}$ if
When all $\mathbb{P}\in\mathcal{P}$ are atomless and $K=1$, then Definitions 7.1 and 3.13 are equivalent to each other and equivalent to Definition 3.1 on approximate p-variables and e-variables. The definitions also agree when $\delta=0$, for instance, $(\varepsilon,0)^{*}$-approximate compound e-variables are identical to $(\varepsilon,0)$-approximate compound e-variables.
In analogy to Definition 3.17, we define (uniformly) *asymptotic compound p-variables and e-variables, requiring, e.g., that $\lim_{m\to\infty}\varepsilon_{m}(\mathbb{P})=\lim_{m\to\infty}\delta_{m}(\mathbb{P})=0$ for each $\mathbb{P}\in\mathcal{P}$. We do not define strongly *asymptotic compound p-variables (or e-variables) since the natural definition is equivalent (under the atomless property) to strongly asymptotic compound p-variables (or e-variables).
Definition 7.1 is implied by Definition 3.13, as can be seen most clearly for approximate compound e-variables by comparing to the result of Proposition 3.15. Moreover, Proposition 3.14 has the following analogous result (whose proof we omit).
Proposition 7.2.
Suppose that for all $k\in\mathcal{K}$, $P_{k}$ (resp. $E_{k}$) is an $(\varepsilon_{k},\delta_{k})$-approximate p-variable (resp. e-variable) for $\mathcal{P}_{k}$. Then, $P_{1},\dots,P_{K}$ (resp. $E_{1},\dots,E_{K}$) are $(\varepsilon,\delta)^{*}$-approximate compound p-variables (resp. e-variables) for the choice:
$\varepsilon(\mathbb{P})=\frac{1}{K}\sum_{k:\mathbb{P}\in\mathcal{P}_{k}}\varepsilon_{k}(\mathbb{P}),\qquad\delta(\mathbb{P})=\frac{1}{K}\sum_{k:\mathbb{P}\in\mathcal{P}_{k}}\delta_{k}(\mathbb{P}).$
We can calibrate *approximate compound p-variables to *approximate compound e-variables.
Proposition 7.3.
Let $\varepsilon:\mathcal{P}\to\mathbb{R}_{+}$ and $\delta:\mathcal{P}\to[0,1]$. Calibrating $(\varepsilon,\delta)^{*}$-approximate compound p-variables yields $(\varepsilon,\delta)^{*}$-approximate compound e-variables, and vice versa.
Theorem 4.4 demonstrated that approximate compound e-variables are naturally compatible with approximate FDR control when using the e-BH procedure. An analogous statement is not true for approximate compound e-variables without further conditions. However it is true in a regime where e-BH makes sufficiently many discoveries. The statement of the following proposition is inspired by Li and Barber (2019, Theorem 2), whose FDR bound also depends on the tail probability of the number of rejections.
Proposition 7.4.
Let $\mathcal{D}$ be the e-BH procedure at level $\alpha\in(0,1)$ and suppose that $E_{1},\ldots,E_{K}$ are $(\varepsilon,\delta)^{*}$-approximate e-variables for $(\mathcal{P}_{1},\ldots,\mathcal{P}_{K})$ under $\mathcal{P}$. Then
$\mathrm{FDR}_{\mathcal{D}}^{\mathbb{P}}\leq\alpha(1+\varepsilon_{\mathbb{P}})+\inf_{\kappa\in(0,1)}\left\{\frac{\delta_{\mathbb{P}}}{\kappa}+\mathbb{P}(R_{\mathcal{D}}<K\kappa)\right\}\leq\alpha(1+\varepsilon_{\mathbb{P}})+\sqrt{\delta_{\mathbb{P}}}+\mathbb{P}(R_{\mathcal{D}}<K\sqrt{\delta_{\mathbb{P}}}).$
Under an asymptotic regime with $\delta_{\mathbb{P}},\varepsilon_{\mathbb{P}}\to 0$ and $\mathbb{P}(R_{\mathcal{D}}<K\kappa)\to 0$ for fixed $\kappa\in(0,1)$, then e-BH asymptotically controls the $\mathrm{FDR}$ at level $\alpha$.